# A Unified Implicit Attention Formulation for Gated-Linear Recurrent Sequence Models

### v1

@@ 3.3 Formulation of RWKV via Attention Matrices
$r_{t}=W_{r}\cdot(u_{r}\otimes x_{t}+(1-u_{r})\otimes x_{t-1}),\quad k_{t}=W_{k}\cdot(u_{k}\otimes x_{t}+(1-u_{k})\otimes x_{t-1},\quad v_{t}=x_{t}$ (16)
@@
$o=\textbf{diag}(\sigma(r))\hat{\alpha}x$ (19)
@@ 3.4 Shared properties
The proposed formulation for Griffin, Mamba, and RWKV is based on the similarities in the structure of the architecture. Our formulation focuses on three main components: (i) the core of the linear attention mechanism (S6 for Mamba, RG-LRU for Griffin, or the WKV operator for RWKV), (ii) a short filter operation implemented via Conv1D in Griffin and Mamba and token shift in RWKV, and (iii) the gate branch, as illustrated in Fig. 1. Additionally, our formulation builds on the following key components: (i) rearranging linear layers and omitting operators that don’t influence the mixer components, (ii) representing the gate branch as a data control linear operator using diagonal matrices instead of vectors, (iii) unrolling the linear recurrent layer to obtain a token-to-token map, and (iv) fusing several cascaded linear operators.
Efficient computation
Naive computation of hidden attention maps is computationally expensive. For each channel, it involves computing an attention matrix, which requires quadratic memory and results in $O(HL^{3})$ time complexity per sample. However, in many applications, we are interested in an attention vector that describes the influence of each token on the CLS token. In such cases, the attention vector map per channel can be efficiently computed using the following three strategies: (i) Compute only a single vector, corresponding to a single row in the attention matrices. (ii) Instead of incorporating dependencies through matrix multiplication by M, we convolve the causal filter vector (with suitable padding) with the CLS-attention vector. Finally, (iii) we avoid materializing the entire gate branch, using only the elements associated with the CLS token. Consequently, we can compute the CLS attention vector with linear memory complexity.
@@ 4 Experiments
We conduct experiments using Mamba, RWKV, and Griffin. Our focus centers on examining the semantic diversity of these models and their utility within explainable AI frameworks.
Visualization of Attention Matrices
In Figure 2, we present a comparative visualization of the attention matrices from Mamba, RWKV, Griffin, and Transformer models. To enhance clarity, we applied the Softmax function to each row of the attention matrices from the transformers and conducted min-max normalization on the absolute values of the matrices from the non-transformer models. In every instance, we used a uniform prompt of size 32. For each architecture, we examined the attention matrices derived from the standard pre-trained models available in the Hugging Face library, including the Recurrent Gemma-2B, RWKV-430M trained on the Pile, and a Mamba-based LLM with 2.8B parameters also trained on the Pile.
As illustrated, the implicit attention matrices of Mamba, Griffin, and RWKV exhibit similarities to those derived from traditional transformers. Echoing findings from [2], we note that dependencies between distant tokens become more apparent in the deeper layers of the model, as shown in the lower rows. Additionally, the attention matrices from RWKV are characterized by distinct horizontal tiles, whereas those from Mamba display a more continuous structure.
Visualization of Attention Maps
Additionally, we present attention maps of the [CLS] token for the ViM model across samples from the ImageNet validation set for various methods, including the previously proposed attention formulation by [2]. for Mamba. We explore three explanation methods: raw attention, attention rollout [1], and attribution following [2, 10], along with a comparison to the ViT counterparts, see Fig. 3. Evidently, the explanation methods that are based on our attention formulation depict much more accurate and sharp maps compared to those of [2] and the ViT counterparts. In Fig. 4 we show similar visualizations in the NLP domain. This indicates that the segmentation maps obtained from our method can be effectively utilized in tasks such as weakly-supervised semantic segmentation, as demonstrated in Appendix. B. More qualitative results for the NLP domain can be found in Appendix. D.
@@
Figure 3: Qualitative results for the different explanation methods for the ViT-small and the Mamba-small models. (a) the original image, (b) the Raw-Attention of [2] over Mamba-Small, (c) Attention-Rollout [2] over Mamba-Small, (d) the Mamba-Attribution [2] method over Mamba-Small, (e) the Raw-Attention with our proposed attention over Mamba-Small, (f) Attention-Rollout with our proposed attention over Mamba-Small, (g) the Mamba-Attribution with our attention over Mamba-Small, (h) Raw-Attention of ViT-Small, (i) Attention Rollout [1] for ViT-Small, (j) Transformer-Attribution [11] for ViT-Small
@@
Figure 4: Qualitative results for NLP, samples are taken from IMDB movie sentiment classification. In (a), we show the results for the previously proposed Mamba’s attention [2], (b) our proposed Mamba’s attention, and in (c) we show our proposed method over RWKV. In the upper row, we show a negative sentiment, and in the lower row, we show a positive sentiment.
Segmentation Tests
We evaluated our proposed Mamba’s implicit attention mechanism by comparing its generated foreground segmentation maps against ground truth from the ImageNet-Segmentation dataset [23]. We employed established metrics (pixel accuracy, mean Intersection-over-Union (mIoU), and mean Average Precision (mAP)) aligning with prior works [10, 11, 38, 25]. Notably, we compared our method with both the ViT and the previously proposed Mamba’s implicit attention from [2].
Results presented in 1 demonstrate that our proposed Mamba’s implicit attention outperforms both the Vision Transformer and the previous proposed Mamba’s attention of  [2] on all metrics over the three different XAI methods. This superior performance suggests the potential of these maps for downstreaming tasks such as weakly supervised semantic segmentation, and mitigating background bias in classifiers [12].
Table 1: Segmentation performance on the ImageNet-Segmentation dataset (percent). Higher is better. The upper part depicts results for ViM-small, and the lower part for ViT-small.
Model Method pixel accuracy mAP mIoU
DeiT S Raw-Attention 59.69 77.25 36.94
VisionMamba-S Raw-Attention [2] 67.64 74.88 45.09
Vision Mamba-S Raw-Attention Ours 67.66 80.00 47.28
DeiT S Attn-Rollout [1] 66.84 80.34 47.85
Vision Mamba-S Attn-Rollout [2] 71.01 80.78 51.51
Vision Mamba-S Attn-Rollout Ours 76.40 83.90 58.48
DeiT S Transformer-Attr [11] 79.26 84.85 60.63
Vision Mamba-S Mamba-Attr [2] 74.72 81.70 54.24
Vision Mamba-S Mamba-Attr Ours 79.60 86.40 62.51
Perturbation Tests
@@
To assess the faithfulness of explanations, we adopted an input perturbation scheme similar to [11, 10]. This method involves systematically masking image pixels based on their predicted relevance from the explanation method. We conducted two experiments: positive and negative perturbation. (1) Positive Perturbation: in this setup, a good explanation prioritizes relevant pixels. We expect the model’s accuracy (specifically, top-1 accuracy) to gradually decrease as we mask pixels in descending order of relevance (most relevant first). (2) Negative Perturbation: Conversely, a robust explanation should maintain model accuracy even when irrelevant pixels are masked. Here, we mask pixels in ascending order of relevance (least relevant first). In both scenarios, we evaluate the explanation quality using the Area-Under-Curve (AUC) metric. AUC considers the model’s accuracy as a function of the percentage of masked pixels (ranging from $10\%$ to $90\%$).
@@
The perturbations results are summarized in Table 2 for various explanation methods under both positive and negative perturbation scenarios on the ImageNet validation set. In the positive perturbation scenario, where lower AUC values indicate better performance, our proposed Mamba’s attention method consistently outperforms the other methods. Specifically, our method achieves the lowest AUC values across all explanation methods, with an AUC of $13.264$ for Raw-Attention, $12.830$ for Attn-Rollout, and a notably low $11.350$ for Attribution. In the negative perturbation scenario, where higher AUC values are better, our method shows the best performance, with AUC values of $47.705$ for Raw-Attention, 50.035 for Attn-Rollout, and $51.310$ for Attribution, outperforming both the method of [2] and the counterpart XAI methods for Vision Transformer.
Perturbation experiments for the NLP domain can be found in Appendix C. In these settings, we fine-tune various LLMs with an additional linear layer classifier and append the [CLS] token to all samples to generate explanation maps, similar to methods used in vision models. These results also demonstrate that our attention formulation surpasses the previous formulation proposed by [2] over the mamba model and is effective also for the RWKV model.
[…]
Positive Perturbation Negative Perturbation
[…]
@@
Ablation study
The architectures we explored implicitly parametrize attention matrices through a composition of several different sub-layers, see Eq.9, 19, and 15. Examples of these sub-layers include linear recurrent layers, gate mechanisms, and other components, such as token-shift or depth-wise convolutions. To better understand the contribution of each of these components, we conduct a sequence of ablation studies. Initially, in Fig.5, we visualize the implicit attention of Mamba, ablating the Conv1D or the gate branch, or focusing solely on the S6 layer. As expected, it seems that the Conv1D layer causes a smoothing effect, and the final Mamba implicit attention is significantly sharper than that of the S6 matrices.
Mamba Mamba w.o Conv Mamba w.o Gate S6
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Comparative visualization of ablated hidden matrices
In Tab.3, we compare several ablation variants of our method. As can be seen, our method, which utilizes all the components of Mamba, achieves a much better score than the ablated versions, illustrating the importance of all components. This experiment reveals that including the Conv1D layer is crucial for high performance. However, the activation has a relatively low impact on the implicit representation. A similar ablation study was conducted for RWKV and presented in Appendix C. It is evident that including the gate branch, as presented in our method, consistently improves performance.
@@
Table 3: Ablation studies for our method on the ImageNet-Segmentation dataset. Higher is better.
Model Method pixel accuracy mAP mIoU
vMamba S Mamba-Attr V2 79.60 86.40 62.51
vMamba S Mamba-Attr V2 w/o act 79.32 86.22 62.41
vMamba S Mamba-Attr V2 w/o conv 70.01 78.87 50.64
vMamba S Mamba-Attr V2 w/o gate 75.11 80.12 55.78
vMamba S Mamba-Attr V2 S6-Attn 72.39 80.09 53.19
@@ 5 Conclusions
Looking ahead, we aim to incorporate additional layers, such as Hyena [44], and HGRN2 [46] into our framework, including their vision-specific variants [16, 18, 66, 51]. Furthermore, we plan to examine how differences in these architectures are reflected in their self-attention matrices and explore whether such insights can reveal more about the inductive biases inherent in each architecture.
@@ Appendix A Representing additional architectures via implicit attention
In sec. 3 we present the formulation of Griffin, RWKV, and Mamba via attention matrices. In this section, we extend our method to other layers, such as RetNet [52] and HGRN [47].
@@
To refine this formulation, we can represent the element-wise multiplication as a matrix multiplication using a diagonal matrix $G=\textbf{diag}(\textbf{swish}(XW_{G}))$. Additionally, per-head statistics can be integrated into $G$. Given that the parallel representation of Retention can be depicted via an attention matrix $R$ (see Eq. 5 in [52]), the entire MSR block simplifies to:
@@
Note that the recurrent rule of the HGRN layer can be computed via an implicit attention represented by a matrix $\alpha_{r}$ (see Eq. 5 in [47]), as follows:
@@
Note that the recurrent rule of the HRGU layer can be computed via an implicit attention represented by a matrix ${\alpha}_{r}$ (see Eq. 5 in [47]), as follows:

### v2

@@ 3.3 Formulation of RWKV via Attention Matrices
$r_{t}=W_{r}\cdot(u_{r}\otimes x_{t}+(1-u_{r})\otimes x_{t-1}),\quad k_{t}=W_{k}\cdot(u_{k}\otimes x_{t}+(1-u_{k})\otimes x_{t-1}),\quad v_{t}=x_{t}$ (18)
@@
$o=\text{diag}(\sigma(r))\hat{\alpha}x$ (21)
@@ 3.4 Shared properties
The proposed formulation for Griffin, Mamba, and RWKV is based on the similarities in the structure of the architecture. Our formulation focuses on three main components: (i) the core of the linear attention mechanism (S6 for Mamba, RG-LRU for Griffin, or the WKV operator for RWKV), (ii) a short filter operation implemented via Conv1D in Griffin and Mamba and token shift in RWKV, and (iii) the gate branch, as illustrated in Fig. 1. Additionally, our formulation builds on the following key components: (1) rearranging linear layers and omitting operators that don’t influence the mixer components, (2) representing the gate branch, activations, and normalization layers as a data control linear operator via diagonal matrices, (3) unrolling the linear recurrent layer to obtain a token-to-token map, and (4) fusing several cascaded linear operators and ignore biases.
@@ 4 Experiments
To assess the effectiveness of our implicit attention formulation, we perform a comprehensive set of experiments. In Sec. 4.1, we begin by visualizing the implicit attention matrices and the corresponding explainability maps built upon them. In Sec 4.2, we demonstrate that integrating our improved attention matrices into existing attribution methods results in SoTA interperablity techniques. We further conduct ablations to analyze the contribution of each architectural component to the overall representation. Finally, in Sec. 4.3, we show how our formulation enables the transfer of performance-enhancing techniques, originally designed for other architectures, to gated RNNs.
@@
4.1 Visualizations
In Figure 2, we present a comparative visualization of the attention matrices from Mamba, RWKV, Griffin, and Transformer models. To enhance clarity, we applied the Softmax function to each row of the attention matrices from the Transformers and conducted min-max normalization on the absolute values of the matrices from the non-Transformer models. In every instance, we used a uniform prompt of size 32. For each model, we examined the attention matrices derived from the standard pre-trained models available in the Hugging Face, including the Recurrent Gemma-2B, RWKV-430M trained on the Pile, and a Mamba-based LLM with 2.8B parameters also trained on the Pile.
As illustrated, the implicit attention matrices of Mamba, Griffin, and RWKV exhibit similarities to those derived from traditional Transformers. Echoing findings from (Ali et al., 2024), we note that dependencies between distant tokens become more apparent in the deeper layers, as shown in the lower rows. Additionally, the matrices from RWKV are characterized by distinct horizontal tiles, whereas those from Mamba display a more continuous structure.
Visualization of Explainability Maps. Sample explainability maps built on top of our implicit attention formulation are resented in Figure 3. This visualization focuses on the rows of the attention maps that are associated with the [CLS] token, as is traditionally done for interpretability purposes. We explore the attention matrices with three explanation methods: raw attention, attention rollout (Abnar & Zuidema, 2020), and attribution following Ali et al. (2024), along with a comparison to the ViT counterparts. Evidently, the explanation methods that are based on our attention formulation (columns e, f, and g) depict much more accurate and sharp maps compared to those of (Ali et al., 2024) and the ViT counterparts. In Fig. 4 we show similar visualizations in the NLP domain. More qualitative results for the NLP domain can be found in Appendix. D.
@@ 4.1 Visualizations
Figure 3: Qualitative results for the different explanation methods for the ViT and ViM, both of small size. (a) The original image, (b) Raw-Attention over ViM, (c) Attention-Rollout over ViM, (d) Mamba-Attribution over ViM, (e) Raw-Attention with our proposed attention over ViM, (f) Attention-Rollout with our proposed attention over ViM, (g) Mamba-Attribution with our proposed attention over ViM, (h) Raw-Attention of ViT, (i) Attention-Rollout for ViT, (j) Transformer-Attribution for ViT. Results for columns (b), (c), and (d) are based on the method of (Ali et al., 2024), and the ViT results on (i), (j) and (k) rely on Chefer et al. (2021b).
@@
Figure 4: Qualitative results for NLP, samples are taken from IMDB movie sentiment classification. In (a), we show the results for the previously proposed Mamba’s attention (Ali et al., 2024), (b) our proposed Mamba’s attention, and in (c) we show our proposed method over RWKV. In the upper row, we show a negative sentiment, and in the lower row, we show a positive sentiment.
4.2 Implicit Attention-Based Attribution
Although empirical evaluation of attribution methods is challenging, this section demonstrates that off-the-shelf techniques, when built on top of our implicit attention formulation, produce SoTA explainability tools. We provide empirical analysis via perturbation and segmentation tests.
@@ 4.2 Implicit Attention-Based Attribution
Perturbation Tests. To assess the faithfulness of explanations, we adopted an input perturbation scheme similar to (Chefer et al., 2021b; a). This method involves systematically masking image pixels based on their predicted relevance from the explanation method. We conducted experiments with both positive and negative perturbations on both NLP and Vision domains. For positive perturbation, a good explanation prioritizes relevant pixels. We expect the model’s accuracy (specifically, top-1 accuracy) to gradually decrease as we mask pixels in descending order of relevance (most relevant first). As for negative Perturbation, a robust explanation should maintain model accuracy even when irrelevant pixels are masked. Here, we mask pixels in ascending order of relevance (least relevant first). In both scenarios, we evaluate the explanation quality using the Area-Under-Curve (AUC) metric. AUC considers the model’s accuracy as a function of the percentage of masked pixels (ranging from $10\%$ to $90\%$).
@@
The perturbations results for vision models are summarized in Table 1 for various explanation methods under both positive and negative perturbation scenarios on the ImageNet validation set. In the positive perturbation scenario, where lower AUC values indicate better performance, our proposed Mamba’s attention method consistently outperforms the other methods. Specifically, our method achieves the lowest AUC values across all explanation methods, with an AUC of $13.264$ for Raw-Attention, $12.830$ for Attn-Rollout, and a notably low $11.350$ for Attribution. In the negative perturbation scenario, where higher AUC values are better, our method shows the best performance, with AUC values of $47.705$ for Raw-Attention, 50.035 for Attn-Rollout, and $51.310$ for Attribution, outperforming both the method of (Ali et al., 2024) and the counterpart XAI methods for ViT.
Table 1: Perturbation Tests for Vision. We present the AUC results (percentages) for the predicted class on the ImageNet validation set. For positive perturbation lower is better, and for negative perturbation higher is better. Previous results by (Ali et al., 2024) denoted by .
Positive Perturbation $\downarrow$ Negative Perturbation $\uparrow$
Mamba Mamba Ours Transformer Mamba Mamba Ours Transformer
@@
In the NLP domain, we conducted perturbation tests in both the zero-shot and fine-tuned settings. In the zero-shot setting, we utilized pre-trained Mamba-based LLMs with sizes of 1.3B and 2.8B on the ARC-E dataset (Clark et al., 2018), which evaluates the reasoning abilities of LLMs. Results are presented in Table 2 and contain activation and pruning perturbations, as described by Ali et al. (2022). It is shown that our explainability method improves upon the baseline of Ali et al. (2024) for both model sizes. Specifically, in the activation scenario, our method improves results by at least 2.2%, and by over 10% in the pruning settings. Similar trends are also evident in the fine-tuned scenario, see Appendix C for these results for both Mamba and RWKV. Taken together, these results demonstrate that our attention formulation is much more precise and better reflects the model’s behavior compared to the formulation proposed by Ali et al. (2024). The same phenomenon occurs with the RWKV model, consistently showing that our formulation can lead to SoTA attribution methods for an entire family of models.
Table 2: Perturbation Tests for NLP. For activation perturbation lower is better, and for pruning perturbation higher is better. Previous results by (Ali et al., 2024) denoted by .
Activation Perturbation (AUAC) $\uparrow$ Pruning Perturbation (AU-MSE)   $\downarrow$
Mamba 1.3B Mamba 1.3B Mamba 2.8B Mamba 2.8B Ours
Attribution of (Ali et al., 2024) 0.915 0.918 1.765 1.239
Our Attribution 0.934 0.939 1.588 1.082
Segmentation Tests. We evaluated our proposed Mamba’s implicit attention mechanism by comparing its generated foreground segmentation maps against ground truth from the ImageNet-Segmentation dataset (Guillaumin et al., 2014). We employed established metrics (pixel accuracy, mean Intersection-over-Union (mIoU), and mean Average Precision (mAP)) aligning with prior works (Chefer et al., 2021b; Nam et al., 2020; Gur et al., 2021). Notably, we compared our method with both the ViT and the previously proposed Mamba’s implicit attention from (Ali et al., 2024).
Results presented in Table 3 demonstrate that our proposed Mamba’s implicit attention outperforms both the ViT and the previous proposed Mamba’s attention of  (Ali et al., 2024) on all metrics over the three different XAI methods. This superior performance suggests the potential of these maps for downstreaming tasks such as weakly supervised semantic segmentation, and mitigating background bias in classifiers (Chefer et al., 2022).
@@
Table 3: Segmentation results on the ImageNet-Segmentation dataset (percent). Higher is better.
Model Method Pixel accuracy $\uparrow$ mAP $\uparrow$ mIoU $\uparrow$
DeiT S Raw-Attention 59.69 77.25 36.94
ViM-S Raw-Attention (Ali et al., 2024) 67.64 74.88 45.09
ViM-S Raw-Attention Ours 67.66 80.00 47.28
DeiT-S Attn-Rollout (Abnar & Zuidema, 2020) 66.84 80.34 47.85
ViM-S Attn-Rollout (Ali et al., 2024) 71.01 80.78 51.51
ViM-S Attn-Rollout Ours 76.40 83.90 58.48
DeiT-S Transformer-Attr (Chefer et al., 2021b) 79.26 84.85 60.63
ViM-S Mamba-Attr (Ali et al., 2024) 74.72 81.70 54.24
ViM-S Mamba-Attr Ours 79.60 86.40 62.51
M. M. w/o Conv M. w/o Gate S6
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
figureComparative visualization of ablated hidden matrices. ’M’ for Mamba.
[…]
tableAblation. ViM-small for ImageNet Segmentation dataset. Higher is better. Method Pixel Acc. mAP mIoU Full Model 79.60 86.40 62.51 w/o SilU 79.32 86.22 62.41 w/o Conv1D 70.01 78.87 50.64 w/o Gate 75.11 80.12 55.78 Only S6 72.39 80.09 53.19
Ablation study. The architectures we explored implicitly parametrize attention matrices through a composition of several different sub-layers, see Eq.9, 21, and 17. Examples of these sub-layers include linear recurrent layers, gate mechanisms, activations, normalization and other components, such as token-shift or depth-wise convolutions. To better understand the contribution of each of these components, we conduct a sequence of ablation studies. Initially, in Fig. 4.2, we visualize the implicit attention of Mamba, ablating the Conv1D or the gate branch, or focusing solely on the S6 layer. As expected, it seems that the Conv1D causes a smoothing effect, and the Mamba matrices are significantly sharper, with more pixels having non-negligible values compared to those of S6.
In Table 4.2, we compare several ablation variants of our method. As can be seen, our method, which utilizes all the components of Mamba, achieves a much better score than the ablated versions, illustrating the importance of all components. This experiment reveals that including the Conv1D and gating mechanism is crucial for high performance and reliable representation. However, the activation has a relatively low impact on these aspects. A similar ablation study was conducted for RWKV and presented in Appendix C, demonstrating similar trends.
4.3 Attribution-Based Performance-Enhancing Techniques
To further demonstrate the practical impact of our representation, we show it can enhances model performance. While various attention-based and explainability-based techniques was previously proposed for improving model performance, our focus is on in-context learning (ICL) and weakly supervised semantic segmentation tasks.
To improve ICL, we adopt the AMPLIFY method of Krishna et al. (2024), a prompt engineering technique designed for few-shot ICL, which leverages post-hoc explanation methods. In our experiments, we use the Mamba-790m model as a proxy, following the same evaluation protocol as AMPLIFY, but with an attribution method that relies on attention matrices. We report the vanilla model performance, and the performance with AMPLIFY with our attribution method and the attribution method of Ali et al. (2024). The results are depicted in Tab. 4, show that our method outperforms the baseline across all tested scenarios except one, with an average margin of 1.2 accuracy points over the amplify baseline, and 9.8% over the vanilla baseline.
Detailed results and experimental settings for weakly supervised semantic segmentation are presented in Appendix B.
Table 4: Performance of various Mamba-based LLMs on Snarks, CommonsenseQA, and Formal Fallacies datasts. We compare the vanilla model performance, and models employ the Amplified method, with our attribution method or the attribution method of Ali et al. (2024) (denoted by ).
Model size Snarks (%) CommonsenseQA (%) Formal Fallacies (%)
1.3B 44.54 52.15 40.13
1.3B + AMPLIFY 53.11 53.55 44.28
1.3B + AMPLIFY (Ours) 56.15 54.72 45.22
2.8B 48.75 53.11 44.67
2.8B + AMPLIFY 58.10 56.10 47.80
2.8B + AMPLIFY (Ours) 60.02 56.08 47.86
@@ 5 Conclusions
Looking ahead, we aim to incorporate additional layers, such as Hyena (Poli et al., 2023), and HGRN2 (Qin et al., 2024a) into our framework, including their vision-specific variants (Duan et al., 2024; Fan et al., 2023; Zimerman & Wolf, 2024; Spravil et al., 2024). Furthermore, we plan to examine how differences in these architectures are reflected in their self-attention matrices and explore whether such insights can reveal more about the inductive bias inherent in each architecture.
@@ Appendix A Representing additional architectures via implicit attention
In sec. 3 we present the formulation of Griffin, RWKV, and Mamba via attention matrices. In this section, we extend our method to other layers, such as RetNet (Sun et al., 2023) and HGRN (Qin et al., 2024b).
@@
To refine this formulation, we can represent the element-wise multiplication as a matrix multiplication using a diagonal matrix $G=\textbf{diag}(\textbf{swish}(XW_{G}))$. Additionally, per-head statistics can be integrated into $G$. Given that the parallel representation of Retention can be depicted via an attention matrix $R$ (see Eq. 5 in (Sun et al., 2023)), the entire MSR block simplifies to:
@@
Note that the recurrent rule of the HGRN layer can be computed via an implicit attention represented by a matrix $\alpha_{r}$ (see Eq. 5 in (Qin et al., 2024b)), as follows:
@@
Note that the recurrent rule of the HRGU layer can be computed via an implicit attention represented by a matrix ${\alpha}_{r}$ (see Eq. 5 in (Qin et al., 2024b)), as follows: